What is the inflation rate?

Consider an economy that produces only chocolate bars. In year 1, the quantity produced is 3 bars and the price is $4. In year 2, the quantity produced is 4 bars andthe price is $5. In year 3, the quantity produced is 5 bars and the price is $6. Year 1 is the base year.
E. What is the inflation rate as measured by the GDP deflator from year 2 to year 3?
Answer #1
y1.....GDP 12$....RGDP 12$

y2....GDP 20$....RGDP 16$.....deflator 125

y3 .....GDP 30$ ....RGDP 20$....deflator 150
answered by: He's not all bad like his reputation </3
Answer #2
nominal GDP is the sum of the quantities*their prices
real GDP is : nominal GDP devided by price index
GDP deflator is : nominal GDP devided by real GDP
so
nominal GDP for year 1 :3*4=12
year 2 and 3 computes like this
for real GDP u must say which kind of index u want to use for example laspiers or ...
hope that these words be useful:)
answered by: jhgjgk
Answer #3
In this case, let us look at the marginal cost in year 1

3 bars=12 dollars
1 bar=4 dollars

year 2
1 bar=5 dollars
Using year 1 as a base we use this formula
(new-base)/base
(5-4)/4=1/4=.25


year 3
1 bar=6 dollars


Using year 1 as a base we use this formula
(new-base)/base
(6-4)/4
=2/4=1/2=.5

Year 2 to year 3 is

.5-.25=.25



There was a 25% increase from year 2 to 3 using a base inflation rate from year 1


Hope this helps
answered by: Ruchee Thapa